how can you recognize a dependent system when solving by addition?
A) look at the graph
B) if the equation you get after substituting turns out to be incorrect, such as 0=9
C) If the identity such as 0=0, results from addition of the equation
D) the lines have inverse slopes
E) Cannot be recognized using this method

Answer by Theo(13342)

Selection C.

If the identity such as 0 - 0 results from addition of the equation.

An example would be identical lines.

x + y = 3
2x + 2y = 6

These lines are identical.

Multiply the first equation by 2 and then subtract it from the second equation to get:

0 + 0 = 0 which becomes 0 = 0.

If you graph these equation, they will show up on your graph as the same line because the equations are identical.